Rosie wants to find the area of a square tabletop .Each side is 20 1/2 inches. What is the area of the tabletop?
Answer by TimothyLamb(4379)

---
a = s*s
s = 20 1/2 = 20.5
a = 20.5*20.5
a = 420.25 in
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php